Green light has a wavelength of 0.00000050 m, and a speed of 300000000 m/s
When the green light goes into glass, it slows down. Its new speed is 200000000m/s
its frequency is 6x10^14 Hz
whats the wavelength in glass?
1
Expert's answer
2020-10-15T10:50:20-0400
"\lambda_{glass}=\frac{v_{glass}}{f}\\\lambda_{glass}=\frac{200000000}{6\cdot 10^{14}}=333\cdot 10^{-9}\ m=333\ nm"